OpenStudy (anonymous):

Solve P = 4(3a – b) for a in terms of P and b. You must show all wor

OpenStudy (anonymous):

P=12a-4b 12a=P+4b a=(P+4b)/12 a=P/12+b/3

OpenStudy (anonymous):

Atchyut is boss